OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) (03.05) Your friend has $100 when he goes to the fair. He spends $10 to enter the fair and $20 on food. Rides at the fair cost $2.00 per ride. Which function can be used to determine how much money he has left over after x rides? f(x) = -2x + 70 f(x) = 2x + 70 f(x) = -2x - 70 f(x) = -30x + $100 -------------------------------------------------------------------------------- Question 2 (Multiple Choice Worth 1 points) (03.05) A school puts on a play. The play costs $1,200 in expenses. The students charge $4.0

OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) (03.05) Your friend has $100 when he goes to the fair. He spends $10 to enter the fair and $20 on food. Rides at the fair cost $2.00 per ride. Which function can be used to determine how much money he has left over after x rides? f(x) = -2x + 70 f(x) = 2x + 70 f(x) = -2x - 70 f(x) = -30x + $100 -------------------------------------------------------------------------------- Question 2 (Multiple Choice Worth 1 points) (03.05) A school puts on a play. The play costs $1,200 in expenses. The students charge $4.00 for tickets. There will be one performance of the play in an auditorium that seats 500 people. What is the domain of the function that shows the profit as a function of the number of tickets sold? The domain is all real numbers from negative 1,200 to positive 2,000 The domain is all integers from negative 1,200 to positive 2,000. The domain is the integers from 0 to 500. The domain is all real numbers from 0 to 500.